OpenStudy (anonymous):

Please help! Will medal!! I need to know how to solve: If s(x) = 2x^2 + 3x - 4, and t(x) = x + 4 then s(x) · t(x) =

OpenStudy (the_fizicx99):

(s*t)(x) ?

OpenStudy (anonymous):

Not a clue. Sorry

OpenStudy (anonymous):

Well, you'd multiply the two values you have. (2x^2 + 3x-4)(x+4) I'd highly suggest factoring the first expression (2x^2 + 3x - 4), if you can. :)

OpenStudy (the_fizicx99):

Hm, well s(x) * t(x) is same as grabbing s of x and t of x and multiplying them... (2x^2 + 3x - 4) * (x + 4) = (2x^2 + 3x - 4)(x + 4, now Foil it. Factoring wouldn't do much.

OpenStudy (the_fizicx99):

Questions? :O

OpenStudy (anonymous):

I don't understand this at all. Sorry

OpenStudy (anonymous):

Did you understand why you multiply the two values? :)

OpenStudy (the_fizicx99):

Hm, suppose that s(x) is 1, and that t(x) is 2, \(\ \sf \color{blue}{s(x)} \times \color{red}{t(x)} = \color{blue}{1} \times \color{red}{2} = 2 \) get it now? :O

OpenStudy (anonymous):

I don't understand how to foil this.

OpenStudy (the_fizicx99):

Except! That because it's more than 1 number or variable, it gets confusing, so use parenthesis. And, s(x) * t(x) = (s*t)(x) ;)

OpenStudy (anonymous):

Well, FOIL works like this: Say you have this: (x+2)(x+3) You'd multiply the first two terms first (x and x) to get x^2. Then you multiply the outside (x and 3) to get 3x Then the inside (2 and x) to get 2x Then the last two terms (2 and 3) to get 6 x^2 + 3x + 2x + 6 x^2+5x+6. FOIL means first, outside, inside, last :)

OpenStudy (anonymous):

I hope that makes sense ^^'

OpenStudy (the_fizicx99):

Same thing! If it helps, switch it around, (x+4)(2x^2 + 3x-4)